News by Date
News by Category
Solar System
White Dwarfs
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Galaxy Clusters
Cosmology/Deep Field
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
Web Shortcuts
Chandra Blog
RSS Feed
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal (opens in new window)
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Operations CXO Status Report

Thursday 23 September 1999 11.00am EST

During the last 24 hours the Observatory passed through the radiation belts and observations of GTO targets continued following belt exit. The current targets include Abell 3667, SS 433, Cygnus X-2 and Capella. These observations are being made with ACIS-I and ACIS-S/HETG. The spacecraft systems continued to operate nominally.

A set of ACIS calibration measurements were made before and after belt passage and for the 3rd orbit in a row showed no further degradation in energy resolution of the front-side illuminated CCD chips.

Return to Status Reports